# Adaptive Step Duration for Precise Foot Placement: Achieving Robust Bipedal Locomotion on Terrains with Restricted Footholds

### v2

@@ I Introduction
Bipedal robots are designed to collaborate with humans while navigating complicated and crowded human environments built for our daily lives. These environments often restrict the robot’s viable footholds, requiring the robot to be able to adaptively plan its gaits in real-time so that the foot always lands in given limited areas. In the literature on bipedal locomotion, such limited areas are often regarded as stepping stones, as illustrated in Fig. 1. Many dynamic walking algorithms fail to successfully tackle stepping stones as they often assume the foot can be placed anywhere within the allowable workspace of the swing foot. A noticeable exception in the literature is the use of full-body trajectory optimization with foothold constraints. For example, a library of gaits can be generated offline with a set of fixed stepping-stone profiles, and then interpolated online given the real-time feedback of the next foot placement area [1] or adapted through an online MPC with a whole-body controller [2]. However, the adaptability of such approaches is limited due to the difficulty in optimizing a full-body trajectory in real-time for any given stepping stone profile.
@@
A popular approach to mitigate the complexity of robot dynamics in gait planning is using reduced-order template models to represent the critical progression of bipedal robots during locomotion. The most commonly used model, the linear inverted pendulum model (LIPM), describes the center of mass (CoM) dynamics with an inverted pendulum with a constant height [3, 4, 5, 6]. To provide a criterion for gait stability of pendulum-based CoM dynamics models, the capturability [7] and the Divergent Component of Motion (DCM) [8] decouple the stable and unstable dynamics based on the viable states (i.e., states that do not lead to falling). Those analyses also show that a variable step duration can increase the range of viable states and allow robots to prevent falling by taking faster and wider steps. However, such a time-varying variable usually leads to nonlinear and non-convex planning problems, which require heavy computing resources and thus impede its implementation in online controllers. Griffin et al. formulate a continuous-time MPC for the time-varying DCM as a mixed-integer quadratically constrained quadratic program (MIQCP) that shows good resistance to force perturbations but suffers from slow online solving (average about 72 ms) [9].
@@
The decoupling of the stable and unstable dynamics through DCM allows the gait planner to regulate both stepping position and timing conveniently. Khadiv et al. introduce the DCM offset to represent the viability bounds of the LIPM model and formulate the desired step position and step duration into an online quadratic programming (QP) [10]. Their foot placement planner finds an optimal foot placement to ensure bounded viable states, thus improving perturbation resistance. There are also extended works for more complicated stepping strategies [11] and more challenging walking perturbation [12], which implies their potential application in bipedal locomotion on stepping stones. Nonetheless, these approaches depend on one-step preview planning, in which the gait features are determined without considering the restrictions on future foothold locations. Consequently, this limits their adaptability to traverse randomly placed stepping stones.
@@
In this work, we develop a multi-step preview gait planning approach to improve the overall stability and long-horizon viability of bipedal walking, even in the presence of severe constraints on the footholds. We introduce a multiple-step Model Predictive Control (MPC) formulation to adjust step durations and stepping positions while walking over randomly located stepping stones. Our formulation uses the modified form of DCM based on the contact angular momentum rather than the linear velocity of the base. This is inspired by the improved performance of the angular momentum-based linear inverted pendulum (ALIP) in bipedal locomotion on various terrains due to better approximation of robot dynamics [13, 14, 15]. We validated the proposed approach with Digit humanoid in simulation under various types of stepping stone profiles and perturbations. Compared with the one-step preview algorithm in  [10], our planner achieves more robust bipedal locomotion over challenging scenarios.
@@
The remainder of the paper is organized as follows: Section II introduces the modified DCM formulation, and Section III proposes a multi-step preview planner that adaptively changes step durations to remain stable on stepping stones. Section IV discusses the modification in low-level task space tracking controller in response to adaptive step durations. Section V shows the simulation results with the Digit robot, and finally Section VI briefly concludes the contributions and discusses future plans.
@@ II Divergent Component of Motion Analysis of Bipedal Locomotion
This section introduces the modified DCM formulation based on the contact angular momentum-based linear inverted pendulum (ALIP) model. We also explicitly discuss the bounds of step positions and durations for stable locomotion.
@@ II-A Center of Mass Dynamics of Bipedal Robots
We consider the ALIP model on a flat ground with an underactuated single-support phase and an instantaneous double-support phase [13]. Using $x_{c},y_{c},z_{c}$ to denote the CoM position in the right-handed contact point frame, the ALIP dynamics are given by
$\begin{split}\dot{x}_{c}&=\frac{L^{y}}{mz_{c}},\quad\quad\ \dot{y}_{c}=-\frac{L^{x}}{mz_{c}},\\ \dot{L}^{x}&=-mgy_{c},\quad\dot{L}^{y}=mgx_{c},\end{split}$ (1)
where $z_{c}$ is the constant CoM height, $L^{x},L^{y}$ are the $x,y$-component of the angular momentum about the contact point (i.e., contact angular momentum), $m$ is the mass of the robot, and $g$ is the gravitational acceleration constant. Compared to the LIPM described in [3], the ALIP model uses the contact angular momentum $L^{x},L^{y}$ instead of the linear velocities of the CoM. This change reduces the mismatch between the template model and the real robot state as it only assumes the centroidal angular momentum is zero instead of the rate of change of the centroidal angular momentum is zero [13].
Figure 2: Schematic view of the DCM evolution with step positions and footprints. During the $k$-th step, the DCM evolves from $\bm{\xi}_{0}^{k}$ to $\bm{\xi}_{T}^{k}$ and is reset to $\bm{\xi}_{0}^{k+1}$ in the new contact frame after the touch-down.
@@ II-B Discrete Dynamics of the initial DCM
The ALIP dynamics in (1) can be transformed into a system that separates the stable and unstable parts of its dynamics, where the latter is its Divergent Component of Motion [8]. Using the states of the ALIP model, the DCM $\bm{\xi}=[\xi_{x},\xi_{y}]^{T}$ in the contact frame is given by
$\bm{\xi}=\begin{bmatrix}x_{c}\\ y_{c}\end{bmatrix}+\frac{1}{\lambda}\frac{1}{mz_{c}}\begin{bmatrix}L^{y}\\ -L^{x}\end{bmatrix},\quad\lambda=\sqrt{\frac{g}{z_{c}}}.$ (2)
@@
where $t$ is the elapsed time (or relative time) from the beginning of the current swing phase, $T$ is the step duration, and $\bm{\xi}_{0}$ is the initial DCM at $t=0$. During a walking step, the DCM $\bm{\xi}(t)$ increase exponentially over the time $t$ and is reset to an initial value after a swing foot touch-down due to the update of the contact frame (see Fig. 2). Suppose that the current step is the $k$-th step ($k\in\mathbb{N^{+}}$), $\bm{\xi}_{0}^{k}$ and $\bm{\xi}_{T}^{k}$ are the initial and final value of the DCM of the $k$-th step, and $\bm{\xi}_{0}^{k+1}$ is the initial value of the ($k$+1)-th step that is also referred to as the DCM offset at the end of the $k$-th step in [10]. Then, the reset map of the DCM is given by
$\begin{split}\bm{\xi}_{0}^{k+1}&=\mathit{\Delta}_{k}(\bm{\xi}_{0}^{k})=\bm{\xi}_{T}^{k}-\mathbf{p}_{TD}^{k}=\bm{\xi}_{0}^{k}e^{\lambda T^{k}}-\mathbf{p}_{TD}^{k},\end{split}$ (4)
where $\mathbf{p}_{TD}^{k}=[x_{TD}^{k},y_{TD}^{k}]^{T}$ is the $k$-th step position in the current contact frame and $T^{k}$ is the $k$-th step duration.
Note that the DCM is innately decoupled in $x$- and $y$-direction. For a 3-D bipedal model, there will be an alternating step width $w_{l/r}$ for the left or right swing foot even if the robot is stepping in place. Thus, we define the lateral step position $y_{TD}^{k}:=w_{l/r}+W^{k}$, where $w_{l/r}$ is a fixed step width of the left or right swing foot during in-place lateral walking, and $W^{k}$ is an offset to $w_{l/r}$ that accounts for the actual lateral movement. That is, given a fixed step width $w>0$, the signed step width for the left swing foot is $w_{l}=w>0$ and for the right $w_{r}=-w<0$. For instance, Fig. 2 shows a rightward walking in the $y$-direction with $W^{k}<0$.
For a periodic gait with a constant step position $\mathbf{p}_{TD}^{k}=[L,w_{l/r}+W]^{T}$ and duration $T^{k}=T$, the nominal value of the initial DCM $\bm{\xi}_{0,\mathrm{nom}}$ can be derived from (4) as
$\bm{\xi}_{0,\mathrm{nom}}=\begin{bmatrix}\xi_{0,x,\mathrm{nom}}\\ \xi_{0,yl/r,\mathrm{nom}}\end{bmatrix}=\begin{bmatrix}\dfrac{L}{e^{\lambda T}-1}\\ \dfrac{w_{l/r}}{e^{\lambda T}+1}+\dfrac{W}{e^{\lambda T}-1}\end{bmatrix},$ (5)
where $\xi_{0,yl/r,\mathrm{nom}}$ of the left or right swing foot are solved by considering two consecutive steps (i.e., one cycle of lateral walking).
In the $x$-direction, given a bounded set $\mathcal{D}$ of allowable step position and $\mathcal{T}$ of allowable step duration, we can obtain a bounded set $\mathcal{X}$ of the nominal initial DCM computed by the values in $\mathcal{D}$ and $\mathcal{T}$ using (5), i.e., $\xi_{0,x,\mathrm{nom}}(L,T):\mathcal{D}\times\mathcal{T}\rightarrow\mathcal{X}$, $\forall L\in\mathcal{D}$, $T\in\mathcal{T}$. Each element in this set represents a particular periodic walking gait corresponding to a foot placement $(L,T)$, which leads to the following proposition.
@@
Given appropriate allowable sets $\mathcal{D}^{k}$, $\mathcal{T}^{k}\subset\mathbb{R}$ of the $k$-th step, for any initial DCM $\xi_{0,x}^{k}$ in $\mathcal{X}^{k}$ that evolves once into $\xi_{0,x}^{k+1}$ using (4), there exists at least one allowable foot placement $(L^{k},T^{k})$ such that $\xi_{0,x}^{k+1}$ still remains in $\mathcal{X}^{k}$.
@@
Let $\mathcal{D}^{k}=[L_{min},L_{max}]$ with $L_{min}<0<L_{max}$ and $\mathcal{T}^{k}=[T_{min},T_{max}]$ with $0<T_{min}<T_{max}$, the bound of $\mathcal{X}^{k}$ is given by (5) as
$\begin{split}\xi_{0,x,min}&=\frac{L_{min}}{e^{\lambda T_{min}}-1}<0,\\ \xi_{0,x,max}&=\frac{L_{max}}{e^{\lambda T_{min}}-1}>0.\end{split}$ (6)
For any $\xi_{0,x}^{k}\in\mathcal{X}^{k}$, the evolved initial DCM by the foot placement $(L^{k},T^{k})$ is $\xi_{0,x}^{k+1}=\xi_{0,x}^{k}e^{\lambda T^{k}}-L^{k}$. Thus, by choosing $(L^{k},T^{k})$ as $(L_{min},T_{min})$ or $(L_{max},T_{min})$, $\xi_{0,x}^{k+1}$ can be bounded as
$\begin{split}\xi_{0,x}^{k+1}&\geq\frac{L_{min}}{e^{\lambda T_{min}}-1}e^{\lambda T_{min}}-L_{min}=\xi_{0,x,min},\\ \xi_{0,x}^{k+1}&\leq\frac{L_{max}}{e^{\lambda T_{min}}-1}e^{\lambda T_{min}}-L_{max}=\xi_{0,x,max},\end{split}$ (7)
i.e., $\xi_{0,x}^{k+1}$ remains in $\mathcal{X}^{k}$. ∎
A similar result also applies to the $y$-direction, where there exist two consecutive foot placements such that any initial DCM $\xi_{0,yl/r}^{k}$ has a bounded evolution. The bound of such $\mathcal{X}$ is referred to as the $\infty$-step capturability bound $d_{\infty}$ (or viability bound) in [7] that theoretically distinguishes the stable walking gaits from the unstable ones, and such a bounded (or viable) evolution of the DCM represents a walking gait that never falls. Moreover, it can be shown that the evolution of the initial DCM in the interior of $\mathcal{X}$ can increase the viability margin, which implies a more ”stable” walking system [16].
For bipedal robotic walking without terrain and physical (e.g., leg length) constraints, the allowable sets of each foot placement are almost identical (though symmetric in the $y$-direction). Thus, we can always find at least one stable gait that converges to a periodic one given an appropriate initial DCM $\bm{\xi}_{0}^{1}$. However, even if the allowable sets of foot placements are constrained, it is still possible to find such stable gaits. This intuitively inspires a foot placement strategy on restricted footholds.

### v3

# Adaptive Step Duration for Accurate Foot Placement: Achieving Robust Bipedal Locomotion on Terrains with Restricted Footholds

@@ Abstract
footnotetext: This work was accepted and presented at the IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS 2025). DOI: 10.1109/IROS60139.2025.11246171.
@@ I Introduction
Bipedal robots are designed to collaborate with humans while navigating complicated and crowded environments built for human use. These environments often restrict the robot’s viable footholds, requiring the robot to adaptively plan its gaits in real-time to ensure each footstep lands within feasible regions. In the literature on bipedal locomotion, such limited regions are often regarded as stepping stones, as illustrated in Fig. 1. Many dynamic walking algorithms fail to successfully tackle stepping stones as they often assume the foot can be placed anywhere within the allowable workspace of the swing foot. A noticeable exception in the literature is the use of full-body trajectory optimization with foothold constraints. For example, a library of gaits is often generated offline with a set of fixed stepping-stone profiles, and then interpolated online given the real-time feedback of the next steppable area [12] or adapted through an online MPC with a whole-body controller [11]. However, the adaptability of such approaches is limited due to the difficulty in optimizing a full-body trajectory in real-time for any given stepping stone profile.
@@
A popular approach to mitigate the complexity of robot dynamics in gait planning is to use reduced-order template models to represent the critical progression of bipedal robots during locomotion. The most commonly used model, the linear inverted pendulum model (LIPM), describes the center of mass (CoM) dynamics with an inverted pendulum with a constant height [7, 17, 14, 13]. To provide a criterion for gait stability of pendulum-based CoM dynamics models, the capturability [10] or the Divergent Component of Motion (DCM) [2] decouples the stable and unstable dynamics based on the viable states (i.e., states that do not lead to falling). These analyses also show that a variable step duration can increase the range of viable states and allow robots to prevent falling by taking faster and wider steps. However, such a time-varying variable usually leads to nonlinear and non-convex planning problems, which require heavy computing resources and thus impede its implementation in online controllers. Griffin et al. formulate a continuous-time MPC for the time-varying DCM as a mixed-integer quadratically constrained quadratic program (MIQCP) that shows good resistance to force perturbations but suffers from slow online solving (average about 72 ms) [6].
@@
The decoupling of the stable and unstable dynamics through DCM allows the gait planner to regulate both stepping position and timing conveniently. Khadiv et al. introduce the DCM offset to represent the viability bounds of the LIPM model and formulate the desired step position and step duration into an online quadratic programming (QP) [8]. Their foot placement planner finds an optimal foot placement to ensure bounded viable states, thus improving perturbation resistance. There are also extended works for more complicated stepping strategies [19] and more challenging walking perturbations [9], which implies their potential application in bipedal locomotion on stepping stones. Nonetheless, these approaches rely on one-step preview planning, in which the gait features are determined without considering the restrictions on future foothold locations. Consequently, this limits their adaptability to traverse randomly placed stepping stones.
@@
In this work, we develop a multi-step preview gait planning approach to improve the overall stability and long-horizon viability of bipedal walking, even in the presence of severe constraints on the footholds. We introduce a multiple-step discrete Model Predictive Control (MPC) formulation to adjust step durations and stepping positions while walking over randomly located stepping stones. Our formulation uses the modified form of the DCM based on the contact angular momentum rather than the linear velocity of the base. This is inspired by the improved performance of the angular momentum-based linear inverted pendulum (ALIP) in bipedal locomotion on various terrains due to better approximation of robot dynamics [5, 4, 3]. We validated the proposed approach with the Digit humanoid in simulation under various types of stepping stone profiles and perturbations. Compared with the one-step preview algorithm in  [8], our planner achieves more robust bipedal locomotion over challenging scenarios.
@@
The remainder of the paper is organized as follows: Section II introduces the modified DCM formulation, and Section III proposes a multi-step preview planner that adaptively changes step durations to remain stable on stepping stones. Section IV discusses the modification in the low-level task space tracking controller in response to adaptive step durations. Section V presents the simulation results using the Digit robot, and Section VI provides a brief conclusion, summarizing the contributions and discussing future plans.
@@ II Divergent Component of Motion Analysis of Bipedal Locomotion
This section introduces the modified DCM formulation based on the contact angular momentum-based linear inverted pendulum (ALIP) model [5]. We also explicitly discuss the bounds of step positions and durations for stable locomotion.
@@ II-A Center of Mass Dynamics of Bipedal Robots
We consider the ALIP model on flat ground with an underactuated single-support phase and an instantaneous double-support phase [5]. Let $(x_{c},y_{c},z_{c})$ be the CoM position in the right-handed contact foot frame, the ALIP dynamics are given by [5]:
$\begin{split}\dot{x}_{c}&=\frac{\mathbf{L}^{y}}{mz_{c}},\quad\quad\ \dot{y}_{c}=-\frac{\mathbf{L}^{x}}{mz_{c}},\\ \dot{\mathbf{L}}^{x}&=-mgy_{c},\quad\dot{\mathbf{L}}^{y}=mgx_{c},\end{split}$ (1)
where $z_{c}$ is the constant CoM height, $\mathbf{L}^{x},\mathbf{L}^{y}$ are the $x,y$-component of the angular momentum about the contact point (i.e., contact angular momentum), $m$ is the mass of the robot, and $g$ is the gravitational acceleration. Compared to the LIPM described in [7], the ALIP model uses the contact angular momentum $\mathbf{L}^{x},\mathbf{L}^{y}$ instead of linear velocities of the CoM. This modification reduces the mismatch between the template model and the real robot state, as it only assumes that the centroidal angular momentum oscillates about zero, rather than its rate of change being zero [5].
Figure 2: Schematic view of the DCM evolution with step positions and footprints. The blue arrows mark the initial DCM at the beginning of steps and the orange ones mark the final DCM. During the $k$-th step, the DCM evolves from $\bm{\xi}_{0,k}$ to $\bm{\xi}_{T,k}$ and is reset to $\bm{\xi}_{0,k+1}$ in the new contact frame after the touchdown.
@@ II-B Discrete Dynamics of the initial DCM
The ALIP dynamics in (1) can be transformed into a form that decouples the stable and unstable components, with the latter referred to as the Divergent Component of Motion (DCM) [2]. Using the states of the ALIP model, the DCM $\bm{\xi}=[\xi^{x},\xi^{y}]^{T}$ in the contact frame is given by
$\bm{\xi}=\begin{bmatrix}x_{c}\\ y_{c}\end{bmatrix}+\frac{1}{\lambda}\frac{1}{mz_{c}}\begin{bmatrix}\mathbf{L}^{y}\\ -\mathbf{L}^{x}\end{bmatrix},\quad\lambda=\sqrt{\frac{g}{z_{c}}}.$ (2)
where the normalized angular momentum terms replace velocity terms from LIPM.
@@
where $t$ is the elapsed time (or relative time) from the beginning of the swing phase, $T$ is the step duration, and $\bm{\xi}_{0}$ is the initial DCM at $t=0$. During a walking step, the DCM $\bm{\xi}(t)$ increases exponentially over the time $t$, and is reset to another initial value after a swing foot touchdown due to the update of the contact frame (see Fig. 2). For the $k$-th step ($k\in\mathbb{N^{+}}$), let $\bm{\xi}_{0,k}$ and $\bm{\xi}_{T,k}$ be the initial and final values of the DCM during the swing phase, and $\bm{\xi}_{0,k+1}$ be the initial value of the (kk+1)-th step which is also referred to as the DCM offset at the end of the $k$-th step in [8]. Then, the reset map of the initial DCM is given by
$\begin{split}\bm{\xi}_{0,k+1}&=\mathit{\Delta}_{k}(\bm{\xi}_{0,k})\\ &=\bm{\xi}_{T,k}-\mathbf{p}_{TD,k}\\ &=\bm{\xi}_{0,k}e^{\lambda T_{k}}-\mathbf{p}_{TD,k},\end{split}$ (4)
where $\mathbf{p}_{TD,k}=[p^{x}_{TD,k},p^{y}_{TD,k}]^{T}$ is the $k$-th step position (i.e., Touch-Down position) in the current contact frame and $T_{k}$ is the $k$-th step duration.
Note that the DCM is innately decoupled in $x$- and $y$-direction. For a 3-D bipedal model, there will be an alternating step width $w^{l/r}$ for the left or right swing foot, even if the robot is stepping in place. Thus, we define the lateral step position $p^{y}_{TD,k}:=w^{l/r}+W_{k}$, where $w^{l/r}$ is a fixed step width of the left or right swing foot during in-place stepping, and $W_{k}$ is an offset to $w^{l/r}$ that accounts for the actual lateral movement. That is, given a fixed value $w>0$, the signed step width for the left swing foot is $w^{l}=w>0$ and for the right $w^{r}=-w<0$. For instance, Fig. 2 shows rightward walking in the $y$-direction with $W_{k}<0$.
For a periodic gait with a constant step position $\mathbf{p}_{TD,k}=[L,w^{l/r}+W]^{T}$ and duration $T_{k}=T$, the nominal value of the initial DCM $\bm{\xi}_{0,\mathrm{nom}}$ can be derived from (4) as
$\bm{\xi}_{0,\mathrm{nom}}=\begin{bmatrix}\xi^{x}_{0,\mathrm{nom}}\\ \xi^{yl/r}_{0,\mathrm{nom}}\end{bmatrix}=\begin{bmatrix}\dfrac{L}{e^{\lambda T}-1}\\ \dfrac{w^{l/r}}{e^{\lambda T}+1}+\dfrac{W}{e^{\lambda T}-1}\end{bmatrix}$, (5)
where $\xi^{yl/r}_{0,\mathrm{nom}}$ of the left or right swing foot are solved by considering two consecutive steps (i.e., one cycle of lateral walking).
In the $x$-direction, given a bounded set $\mathcal{D}$ of allowable step position and $\mathcal{T}$ of allowable step duration, we can obtain a bounded set $\mathcal{X}$ of the nominal initial DCM computed by the values in $\mathcal{D}$ and $\mathcal{T}$ using (5), i.e., $\xi^{x}_{0,\mathrm{nom}}(L,T):\mathcal{D}\times\mathcal{T}\rightarrow\mathcal{X}$, $\forall L\in\mathcal{D}$, $T\in\mathcal{T}$. Each element in this set represents a particular periodic walking gait corresponding to a foot placement $(L,T)$, which leads to the following proposition.
@@
Given appropriate allowable sets $\mathcal{D}$, $\mathcal{T}\subset\mathbb{R}$ of the $k$-th step, for any initial DCM $\xi^{x}_{0,k}$ in the corresponding $\mathcal{X}$ that evolves once into $\xi^{x}_{0,k+1}$ using (4), there exists at least one allowable foot placement $(L_{k},T_{k})$ such that $\xi^{x}_{0,k+1}$ still remains in $\mathcal{X}$.
@@
Let $\mathcal{D}=[L_{\mathrm{min}},L_{\mathrm{max}}]$ with $L_{\mathrm{min}}<0<L_{\mathrm{max}}$ and $\mathcal{T}=[T_{\mathrm{min}},T_{\mathrm{max}}]$ with $0<T_{\mathrm{min}}<T_{\mathrm{max}}$, the corresponding $\mathcal{X}=[\xi^{x}_{0,\mathrm{min}},\xi^{x}_{0,\mathrm{max}}]$ is given by (5) as
$\begin{split}\xi^{x}_{0,\mathrm{min}}&=\frac{L_{\mathrm{min}}}{e^{\lambda T_{\mathrm{min}}}-1}<0,\\ \xi^{x}_{0,\mathrm{max}}&=\frac{L_{\mathrm{max}}}{e^{\lambda T_{\mathrm{min}}}-1}>0.\end{split}$ (6)
For any $\xi^{x}_{0,k}\in\mathcal{X}$, the evolved initial DCM by the foot placement $(L_{k},T_{k})$ is $\xi^{x}_{0,k+1}=\xi^{x}_{0,k}e^{\lambda T_{k}}-L_{k}$. Then, we can choose $(L_{k},T_{k})$ as boundary values $(L_{\mathrm{min}},T_{\mathrm{min}})$ or $(L_{\mathrm{max}},T_{\mathrm{min}})$, and bound $\xi^{x}_{0,k+1}$ as follows:
$\begin{split}\xi^{x}_{0,k+1}&\geq(\frac{L_{\mathrm{min}}}{e^{\lambda T_{\mathrm{min}}}-1})e^{\lambda T_{\mathrm{min}}}-L_{\mathrm{min}}=\xi^{x}_{0,\mathrm{min}},\\ \xi^{x}_{0,k+1}&\leq(\frac{L_{\mathrm{max}}}{e^{\lambda T_{\mathrm{min}}}-1})e^{\lambda T_{\mathrm{min}}}-L_{\mathrm{max}}=\xi^{x}_{0,\mathrm{max}},\end{split}$ (7)
i.e., $\xi^{x}_{0,k+1}$ remains in $\mathcal{X}$. ∎
A similar result also applies to the $y$-direction, where there exist two consecutive foot placements such that any initial DCM $\xi^{yl/r}_{0,k}$ has a bounded evolution. The bound of such $\mathcal{X}$ is referred to as the $\infty$-step capturability bound $d_{\infty}$ (or viability bound) in [10] that theoretically distinguishes the stable walking gaits from the unstable ones, and such a bounded (i.e., viable) evolution of the DCM represents a walking gait that can be stabilized. Indeed, it can be shown that any evolution of the initial DCM in the interior of $\mathcal{X}$} can increase the viability margin, which implies an “asymptotically stable” walking system [16].
For bipedal robotic walking on flat ground without terrain constraints, the allowable sets of each foot placement are sufficiently large and almost identical (though symmetric in the $y$-direction), thus we can always find at least one stable gait that converges to a desired periodic one given an appropriate initial DCM $\bm{\xi}_{0,1}$. However, even if the allowable sets of foot placements are constrained and the robot state is perturbed, as long as there exists an evolution of the initial DCM that ensures each foot placement lies within its corresponding allowable set, it is still possible to find a viable though aperiodic gait with the bounded DCM. This intuitively inspires a foot placement strategy on restricted footholds with force perturbations.